Improving IV&V Techniques Through the Analysis of Anomalies

Tim Menzies

Lane Department of Computer Science and Electrical Engineering

West Virginia University

tim@menzies.us
Download from http://unbox.org/wisp/trunk/silap/doc/tb.pdf

August 3, 2007

Problem

Problem Contents Slide 1 Use it or Lose it First section

SOUND BITES

Control (of software processes) is possible; Many examples of successful mining of

NASA data Too many mine collapses (data collection

cancelled); Time to register and us all data resources;

Problem Contents Slide 1 Use it or Lose it First section Slide 2 Slide 3

Contents

Problem Contents Slide 1 Use it or Lose it First section

Problem Contents Slide 1 Use it or Lose it

First section Slide 2 Slide 3

Slide 1

Problem Contents Slide 1 Use it or Lose it

First section

This is the first slide

There is nothing special about it.

Problem Contents Slide 1 Use it or Lose it

First section

This is the first slide

There is nothing special about it.

Introduction: Use it or Lose it (and this is the last “business-level” slide)

Problem Contents Slide 1 Use it or Lose it First section

NASA data = active repository or data tomb?

write once; read never; buried; doomed;

If an organization spends millions of dollars on data collection and archiving...

It should spend tens of thousands (at least) in analyzing that data.

NASA IV&V is in a unique position to review and comment on much of the NASA software enterprise.

We see more, but what have we

learned? What can we tell?

First section

Problem Contents Slide 1 Use it or Lose it

First section Slide 2 Slide 3 Problem Contents Slide 1 Use it or Lose it

First section Slide 2 Slide 3

Here

we

demonstrate

the itemize environment

Problem Contents Slide 1 Use it or Lose it

First section Slide 2 Slide 3

Here

we

demonstrate

the itemize environment

Problem Contents Slide 1 Use it or Lose it

First section Slide 2 Slide 3

Here

we

demonstrate

the itemize environment

Problem Contents Slide 1 Use it or Lose it

First section Slide 2 Slide 3

Here

we

demonstrate

the itemize environment

Problem Contents Slide 1 Use it or Lose it

First section Slide 2 Slide 3

1. Here

(a) we

i. demonstrate

A. the enumerate environment

Problem Contents Slide 1 Use it or Lose it

First section Slide 2 Slide 3

1. Here

(a) we

i. demonstrate

A. the enumerate environment

Problem Contents Slide 1 Use it or Lose it

First section Slide 2 Slide 3

1. Here

(a) we

i. demonstrate

A. the enumerate environment

Problem Contents Slide 1 Use it or Lose it

First section Slide 2 Slide 3

1. Here

(a) we

i. demonstrate

A. the enumerate environment